# Memory-Dominated Quantum Criticality as a Universal Route to High-Temperature Superconductivity

### 09 Mar 2026 v3

@@ Appendix A Appendix A: Relaxation-rate spectra as the fundamental dynamical basis of collective dynamics
$S_{0}=\int dt\,d^{d}r\;\tilde{\phi}\big(\partial_{t}+\Gamma(r-c\nabla^{2})\big)\phi-\int dt\,d^{d}r\,D\,\tilde{\phi}^{2}.$ (S64)
@@
up to the conventional factor $\Gamma$ and the $i\omega$ term in the retarded response.
@@ Appendix C Appendix C: Exact emergence of a memory kernel from integrating out an OU reservoir
$\Big[-i\omega+r+\Sigma^{R}(\omega)\Big]\phi(\omega)=\eta(\omega)+\int_{0}^{\Lambda}d\lambda\,\frac{g(\lambda)}{\lambda-i\omega}\zeta_{\lambda}(\omega),$ (S91)
@@
$\Sigma^{R}(\omega)=\int_{0}^{\Lambda}d\lambda\,\frac{g(\lambda)^{2}}{\lambda-i\omega}=\int_{0}^{\Lambda}d\lambda\,\frac{\rho(\lambda)}{\lambda-i\omega}.$ (S92)
We use the convention $G_{R}^{-1}(\omega)=-i\omega+r+\Sigma^{R}(\omega)$; other sign conventions for the self-energy appear in the literature.
@@
$\mathrm{Re}\,\Sigma^{R}(\omega)$ $=\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)\frac{\lambda}{\lambda^{2}+\omega^{2}},$ (S94)
$\mathrm{Im}\,\Sigma^{R}(\omega)$ $=\omega\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)\frac{1}{\lambda^{2}+\omega^{2}}.$ (S95)
@@
$\mathrm{Im}\,\Sigma^{R}(\omega)$ $\simeq\omega\rho_{0}\int_{0}^{\Lambda}\frac{d\lambda}{\lambda^{2}+\omega^{2}}=\omega\rho_{0}\left[\frac{1}{|\omega|}\arctan\!\Big(\frac{\Lambda}{|\omega|}\Big)\right]$
@@
$\mathrm{Re}\,\Sigma^{R}(\omega)$ $\simeq\rho_{0}\int_{0}^{\Lambda}d\lambda\,\frac{\lambda}{\lambda^{2}+\omega^{2}}=\frac{\rho_{0}}{2}\int_{0}^{\Lambda}d\lambda\,\frac{2\lambda}{\lambda^{2}+\omega^{2}}$
@@
Summary. Equations (S81)–(S85) show that integrating out a continuum of OU relaxation modes yields an exact generalized Langevin equation for the collective coordinate with a causal memory kernel $K(t)$. The kernel is the Laplace transform of the TDOS. A finite TDOS at the origin produces the universal long-memory tail $K(t)\sim 1/t$, and the corresponding retarded self-energy exhibits the marginal nonanalytic structure $\mathrm{Im}\,\Sigma^{R}(\omega)\propto\mathrm{sgn}(\omega)$ together with a logarithmic real part. This makes explicit that memory-dominated infrared dynamics arises from the self-organized accumulation of slow internal relaxation modes, rather than from coupling to a fast Markovian bath.
@@
Appendix D Appendix D: Finite-temperature regulation of the infrared divergence
In this Appendix we explicitly demonstrate how the zero–temperature infrared divergence of the polarization kernel, $\Pi(\omega)\sim 1/|\omega|$, is regulated at finite temperature, yielding the universal scaling $\Pi(T)\sim 1/T$. This establishes rigorously that the thermal scale acts as the infrared cutoff in the memory–dominated critical regime.
(I) Spectral integral representation: In the slow–mode reservoir regime the Cooper-channel polarization admits the spectral form
$\Pi(\omega)=\int_{0}^{\Lambda}d\lambda\;\frac{\rho_{\rm eff}(\lambda)}{\lambda^{2}+\omega^{2}}.$ (S101)
[…]
$\Pi(\omega)\simeq\rho_{0}\int_{0}^{\Lambda}d\lambda\;\frac{1}{\lambda^{2}+\omega^{2}}.$ (S102)
The elementary integral yields
$\int_{0}^{\Lambda}\frac{d\lambda}{\lambda^{2}+\omega^{2}}=\frac{1}{|\omega|}\tan^{-1}\!\left(\frac{\Lambda}{|\omega|}\right).$ (S103)
In the low–frequency limit $|\omega|\ll\Lambda$, one obtains
$\Pi(\omega)\simeq\frac{\pi\rho_{0}}{2}\,\frac{1}{|\omega|},$ (S104)
which expresses the nonanalytic infrared enhancement at zero temperature.
(II) Thermal infrared cutoff: At finite temperature the imaginary–time direction is compactified to the interval $\tau\in[0,\beta]$ with $\beta=1/T$, and dynamical correlations are expressed in terms of discrete Matsubara frequencies. As a consequence, continuous access to arbitrarily small frequencies is lost, and the thermal scale provides an effective infrared regulator.
Operationally, the zero–temperature denominator $\omega^{2}$ in Eq. (S102) is replaced by the minimal dynamical scale of order $T^{2}$, yielding
$\Pi(T)\simeq\rho_{0}\int_{0}^{\Lambda}d\lambda\;\frac{1}{\lambda^{2}+(cT)^{2}},$ (S105)
where $c$ is a numerical constant of order unity determined by the Matsubara spectrum.
Evaluating the integral gives
$\Pi(T)=\rho_{0}\,\frac{1}{cT}\tan^{-1}\!\left(\frac{\Lambda}{cT}\right)\simeq\frac{\pi\rho_{0}}{2c}\,\frac{1}{T},\qquad T\ll\Lambda.$ (S106)
Thus the infrared divergence is universally cut off by the thermal scale,
$\Pi(T)\propto\frac{1}{T}.$ (S107)
(III) Equivalent Matsubara summation: The same result follows directly from the finite–temperature Matsubara representation of the polarization bubble,
$\Pi(0,T)\sim T\sum_{\omega_{m}}\int_{0}^{\Lambda}d\lambda\;\frac{\rho_{0}}{\lambda^{2}+\omega_{m}^{2}},\qquad\omega_{m}=(2m+1)\pi T.$ (S108)
Using the standard identity
$T\sum_{\omega_{m}}\frac{1}{\lambda^{2}+\omega_{m}^{2}}=\frac{1}{2\lambda}\tanh\!\left(\frac{\lambda}{2T}\right),$ (S109)
and expanding for $\lambda\ll T$,
$\tanh\!\left(\frac{\lambda}{2T}\right)\simeq\frac{\lambda}{2T},$ (S110)
@@ Appendix D Appendix D: Finite-temperature regulation of the infrared divergence
$T\sum_{\omega_{m}}\frac{1}{\lambda^{2}+\omega_{m}^{2}}\simeq\frac{1}{2T}.$ (S111)
Substitution into Eq. (S108) yields
$\Pi(0,T)\sim\int_{0}^{\Lambda}d\lambda\;\rho_{0}\frac{1}{2T}\propto\frac{\rho_{0}}{T},$ (S112)
confirming that finite temperature universally regulates the infrared divergence.
(IV) Implication for the superconducting transition: The Thouless criterion,
$\chi_{\Delta}^{-1}(0,T_{c})=0\qquad\Longleftrightarrow\qquad\frac{1}{g_{\rm pair}}=\Pi(T_{c}),$ (S113)
therefore gives
$T_{c}\sim\rho_{0}\,g_{\rm pair},$ (S114)
or equivalently $T_{c}\sim\rho_{0}E_{\rm pair}$ in terms of the bare electronic pairing scale. The exponential suppression characteristic of BCS theory is thus replaced by algebraic scaling controlled by the infrared spectral weight of the slow–mode reservoir.
@@ Appendix F Appendix F: Possible connections between flat TDOS and scale–free dynamical phenomena
In this Appendix we briefly discuss a broader dynamical perspective suggested by the relaxation–spectrum framework developed in the main text. The central quantity of the theory is the density of relaxation rates associated with the Liouvillian generator of the coarse–grained dynamics.
Near dynamical critical regimes, the spectrum of relaxation rates naturally develops a dense continuum extending toward vanishing decay rates. In particular, an important limiting case is a finite infrared density of slow modes,
@@
This condition corresponds to a situation in which relaxation processes exist across arbitrarily long time scales. As a result, the system cannot be characterized by a single relaxation time, and long–time memory effects emerge naturally.
(I) Long–time memory kernel: The dynamical kernel associated with the relaxation spectrum can be written as
@@
If the TDOS approaches a constant for $\lambda\ll\Lambda$, $\rho(\lambda)\approx\rho_{0}$, the long–time behavior becomes
@@
For $t\gg\Lambda^{-1}$ one therefore obtains the algebraic decay
@@
(II) Relation to scale–free relaxation statistics: Distributions of relaxation times with power–law statistics are widely encountered in complex systems, including models of self–organized criticality (SOC). In many SOC models the lifetime distribution of events takes the approximate form
@@
Using the relation between lifetime and relaxation rate $\lambda=1/\tau$, one finds
$\rho(\lambda)=P(\tau)\left|\frac{d\tau}{d\lambda}\right|=P(1/\lambda)\frac{1}{\lambda^{2}}.$ (S136)
Substituting $P(\tau)\sim\tau^{-2}$ yields
$\rho(\lambda)\sim{\rm const}.$ (S137)
Thus a flat infrared TDOS corresponds to a relaxation spectrum that is consistent with scale–free lifetime statistics. This observation suggests a possible dynamical connection between the TDOS framework and the emergence of SOC–type temporal scaling.
(III) Implications for low–frequency noise: A broad distribution of relaxation times is also known to generate 1/f1/f–type noise spectra. Within the present framework the noise spectrum can be written as
$S(\omega)\sim\int d\lambda\,\frac{\rho(\lambda)\lambda}{\lambda^{2}+\omega^{2}}.$ (S138)
When the infrared TDOS approaches a constant, the dominant contribution arises from $\lambda\lesssim\omega$, which yields approximately
@@
This provides a natural dynamical interpretation of the ubiquitous $1/f$ noise observed in many correlated systems.
@@
The resulting spectral function
$A(\omega,q)=\frac{2\Gamma}{\left[\Delta_{q}+\rho_{0}\ln(\Lambda/|\omega|)\right]^{2}+\Gamma^{2}}$ (S141)
describes a broad continuum rather than a sharp quasiparticle peak. Such continua have been reported in several strongly correlated systems, including strange metals, although the precise microscopic origin remains an open question.
@@
For $\rho(\lambda\rightarrow 0)=\rho_{0}$ one finds
@@
which leads to the algebraic transition scale
@@
Summary. These observations suggest that the infrared organization of the relaxation spectrum may provide a common dynamical language for several scale–free phenomena observed in correlated systems. Within this perspective the central object is not a single quasiparticle mode but the entire spectrum of collective relaxation processes. Further investigation of this spectral viewpoint may help clarify the role of dynamical criticality in a wide range of nonequilibrium and strongly correlated systems.

### 16 Mar 2026 v4

@@ Appendix A Appendix A: Relaxation-rate spectra as the fundamental dynamical basis of collective dynamics
$S_{0}=\int dt\,d^{d}r\;\tilde{\phi}\big(\partial_{t}+(r-c\nabla^{2})\big)\phi-\int dt\,d^{d}r\,D\,\tilde{\phi}^{2}.$ (S64)
@@
up to the conventional $i\omega$ term of the retarded response.
@@ Appendix C Appendix C: Exact emergence of a memory kernel from integrating out an OU reservoir
$\Big[-i\omega+r+\Sigma_{R}(\omega)\Big]\phi(\omega)=\eta(\omega)+\int_{0}^{\Lambda}d\lambda\,\frac{g(\lambda)}{\lambda-i\omega}\zeta_{\lambda}(\omega),$ (S91)
@@
$\Sigma_{R}(\omega)=\int_{0}^{\Lambda}d\lambda\,\frac{g(\lambda)^{2}}{\lambda-i\omega}=\int_{0}^{\Lambda}d\lambda\,\frac{\rho(\lambda)}{\lambda-i\omega}.$ (S92)
We use the convention $G_{R}^{-1}(\omega)=-i\omega+r+\Sigma_{R}(\omega)$; other sign conventions for the self-energy appear in the literature.
@@
$\mathrm{Re}\,\Sigma_{R}(\omega)$ $=\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)\frac{\lambda}{\lambda^{2}+\omega^{2}},$ (S94)
$\mathrm{Im}\,\Sigma_{R}(\omega)$ $=\omega\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)\frac{1}{\lambda^{2}+\omega^{2}}.$ (S95)
@@
$\mathrm{Im}\,\Sigma_{R}(\omega)$ $\simeq\omega\rho_{0}\int_{0}^{\Lambda}\frac{d\lambda}{\lambda^{2}+\omega^{2}}=\omega\rho_{0}\left[\frac{1}{|\omega|}\arctan\!\Big(\frac{\Lambda}{|\omega|}\Big)\right]$
@@
$\mathrm{Re}\,\Sigma_{R}(\omega)$ $\simeq\rho_{0}\int_{0}^{\Lambda}d\lambda\,\frac{\lambda}{\lambda^{2}+\omega^{2}}=\frac{\rho_{0}}{2}\int_{0}^{\Lambda}d\lambda\,\frac{2\lambda}{\lambda^{2}+\omega^{2}}$
@@
Summary. Equations (S81)–(S85) show that integrating out a continuum of OU relaxation modes yields an exact generalized Langevin equation for the collective coordinate with a causal memory kernel $K(t)$. The kernel is the Laplace transform of the TDOS. A finite TDOS at the origin produces the universal long-memory tail $K(t)\sim 1/t$, and the corresponding retarded self-energy exhibits the marginal nonanalytic structure $\mathrm{Im}\,\Sigma_{R}(\omega)\propto\mathrm{sgn}(\omega)$ together with a logarithmic real part. This makes explicit that memory-dominated infrared dynamics arises from the self-organized accumulation of slow internal relaxation modes, rather than from coupling to a fast Markovian bath.
@@
Appendix D Appendix D: Derivation of the Cooper-channel polarization kernel from a relaxational mode continuum
In this Appendix we present an explicit derivation of the polarization kernel used in the Cooper channel within the relaxation-spectrum framework. The Cooper-channel object relevant for pairing enhancement is the fluctuation (or correlator) spectrum of the collective pair field.
The construction below should be understood, in the same spirit as Appendix B, as an effective infrared description after microscopic electronic degrees of freedom have already been integrated out. At that stage the relevant dynamical variables are collective pair modes, and the polarization kernel is obtained from their stochastic two-point correlations rather than from an explicit microscopic fermionic loop.
(I) Single relaxational pair mode: Consider a single complex collective pair mode $\Delta_{\lambda}(t)$ with relaxation rate $\lambda>0$ obeying the linear Langevin equation
$\partial_{t}\Delta_{\lambda}(t)=-\lambda\,\Delta_{\lambda}(t)+\xi_{\lambda}(t),$ (S101)
where $\xi_{\lambda}(t)$ is Gaussian white noise with
$\langle\xi_{\lambda}(t)\rangle=0,\qquad\langle\xi_{\lambda}(t)\xi_{\lambda^{\prime}}^{*}(t^{\prime})\rangle=2D_{\lambda}\,\delta(\lambda-\lambda^{\prime})\delta(t-t^{\prime}).$ (S102)
Here $D_{\lambda}$ denotes the noise strength associated with the mode $\lambda$.
In the stationary regime, obtained by sending the initial time to -$\infty$, the exact solution of Eq. (S101) is
$\Delta_{\lambda}(t)=\int_{-\infty}^{t}ds\,e^{-\lambda(t-s)}\xi_{\lambda}(s).$ (S103)
Differentiating Eq. (S103) directly yields
$\partial_{t}\Delta_{\lambda}(t)$ $=\xi_{\lambda}(t)-\lambda\int_{-\infty}^{t}ds\,e^{-\lambda(t-s)}\xi_{\lambda}(s)$
$=\xi_{\lambda}(t)-\lambda\Delta_{\lambda}(t),$ (S104)
which confirms Eq. (S101).
(II) Time-domain two-point correlation function: We now compute the stationary correlator
$C_{\lambda}(\tau)\equiv\langle\Delta_{\lambda}(\tau)\Delta_{\lambda}^{*}(0)\rangle.$ (S105)
For $\tau>0$, inserting Eq. (S103) gives
$C_{\lambda}(\tau)$ $=\int_{-\infty}^{\tau}ds\int_{-\infty}^{0}ds^{\prime}\,e^{-\lambda(\tau-s)}e^{\lambda s^{\prime}}\,\langle\xi_{\lambda}(s)\xi_{\lambda}^{*}(s^{\prime})\rangle$
$=2D_{\lambda}\int_{-\infty}^{\tau}ds\int_{-\infty}^{0}ds^{\prime}\,e^{-\lambda(\tau-s)}e^{\lambda s^{\prime}}\delta(s-s^{\prime}).$ (S106)
Because the delta function enforces $s=s^{\prime}$, while the second integral restricts $s^{\prime}\leq 0$, the effective integration domain reduces to $s\leq 0$. Therefore
$C_{\lambda}(\tau)$ $=2D_{\lambda}\int_{-\infty}^{0}ds\,e^{-\lambda(\tau-s)}e^{\lambda s}$
$=2D_{\lambda}e^{-\lambda\tau}\int_{-\infty}^{0}ds\,e^{2\lambda s}$
$=2D_{\lambda}e^{-\lambda\tau}\left[\frac{e^{2\lambda s}}{2\lambda}\right]_{-\infty}^{0}$
$=\frac{D_{\lambda}}{\lambda}e^{-\lambda\tau}.$ (S107)
By stationarity and Hermitian symmetry, the result extends to all times as
$C_{\lambda}(t)=\frac{D_{\lambda}}{\lambda}\,e^{-\lambda|t|}.$ (S108)
Equation (S108) is the exact two-point correlation function of a single Ornstein–Uhlenbeck pair mode. The factor $1/\lambda$ in front of the exponential is crucial. It is precisely this factor that removes the extra numerator $\lambda$ appearing in the Fourier transform of the bare retarded kernel.
(III) Frequency-domain spectral density: The mode-resolved fluctuation spectrum is defined as
$S_{\lambda}(\omega)\equiv\int_{-\infty}^{\infty}dt\,e^{i\omega t}C_{\lambda}(t).$ (S109)
Substituting Eq. (S108) gives
$S_{\lambda}(\omega)=\frac{D_{\lambda}}{\lambda}\int_{-\infty}^{\infty}dt\,e^{i\omega t}e^{-\lambda|t|}.$ (S110)
Splitting the integral into positive and negative times,
$S_{\lambda}(\omega)$ $=\frac{D_{\lambda}}{\lambda}\left[\int_{0}^{\infty}dt\,e^{-(\lambda-i\omega)t}+\int_{-\infty}^{0}dt\,e^{(\lambda+i\omega)t}\right].$ (S111)
Changing variables u=tu=-t in the second term yields
$\int_{-\infty}^{0}dt\,e^{(\lambda+i\omega)t}=\int_{0}^{\infty}du\,e^{-(\lambda+i\omega)u},$ (S112)
so that
$S_{\lambda}(\omega)=\frac{D_{\lambda}}{\lambda}\left[\int_{0}^{\infty}dt\,e^{-(\lambda-i\omega)t}+\int_{0}^{\infty}dt\,e^{-(\lambda+i\omega)t}\right].$ (S113)
Using
$\int_{0}^{\infty}dt\,e^{-at}=\frac{1}{a},\qquad\Re a>0,$ (S114)
we obtain
$S_{\lambda}(\omega)$ $=\frac{D_{\lambda}}{\lambda}\left(\frac{1}{\lambda-i\omega}+\frac{1}{\lambda+i\omega}\right)$
$=\frac{D_{\lambda}}{\lambda}\frac{2\lambda}{\lambda^{2}+\omega^{2}}.$ (S115)
Hence
$S_{\lambda}(\omega)=\frac{2D_{\lambda}}{\lambda^{2}+\omega^{2}}.$ (S116)
This is the key result. The Fourier transform of the exponential memory kernel itself produces a numerator $\lambda$, but the stationary correlator contains a compensating factor $1/\lambda$. As a result, the exact fluctuation spectrum carries no additional factor of $\lambda$ in the numerator.
(IV) Equivalent derivation in frequency space: The same result follows directly in frequency space. Fourier transforming Eq. (S101) gives
$(-i\omega+\lambda)\Delta_{\lambda}(\omega)=\xi_{\lambda}(\omega),$ (S117)
hence
$\Delta_{\lambda}(\omega)=\frac{1}{\lambda-i\omega}\,\xi_{\lambda}(\omega).$ (S118)
The noise correlator in frequency space is
$\langle\xi_{\lambda}(\omega)\xi_{\lambda^{\prime}}^{*}(\omega^{\prime})\rangle=2\pi\,2D_{\lambda}\,\delta(\lambda-\lambda^{\prime})\delta(\omega-\omega^{\prime}).$ (S119)
Therefore
$\langle\Delta_{\lambda}(\omega)\Delta_{\lambda^{\prime}}^{*}(\omega^{\prime})\rangle$ $=\frac{\langle\xi_{\lambda}(\omega)\xi_{\lambda^{\prime}}^{*}(\omega^{\prime})\rangle}{(\lambda-i\omega)(\lambda^{\prime}+i\omega^{\prime})}$
$=2\pi\,\delta(\lambda-\lambda^{\prime})\delta(\omega-\omega^{\prime})\frac{2D_{\lambda}}{\lambda^{2}+\omega^{2}}.$ (S120)
Defining the spectral density through
$\langle\Delta_{\lambda}(\omega)\Delta_{\lambda}^{*}(\omega^{\prime})\rangle=2\pi\delta(\omega-\omega^{\prime})S_{\lambda}(\omega),$ (S121)
one again recovers Eq. (S116).
This derivation makes clear that the mode-resolved pair correlator is proportional to $D_{\lambda}^{R}(\omega)D_{\lambda}^{A}(\omega)$, not to the real part of a single retarded propagator. Indeed,
$D_{\lambda}^{R}(\omega)=\frac{1}{\lambda-i\omega},\qquad D_{\lambda}^{A}(\omega)=\frac{1}{\lambda+i\omega},$ (S122)
so that
$D_{\lambda}^{R}(\omega)D_{\lambda}^{A}(\omega)=\frac{1}{\lambda^{2}+\omega^{2}}.$ (S123)
(V) From a single mode to the Cooper-channel polarization: After microscopic electronic degrees of freedom are integrated out, the collective Cooper sector is represented as a continuum of relaxational pair modes. Let $\rho_{\rm pair}(\lambda)$ denote the effective density of pair relaxation rates, including both mode degeneracy and noise weight. The full polarization kernel is then obtained by summing the mode-resolved fluctuation spectrum over the continuum,
$\Pi(\omega)=\int_{0}^{\Lambda}d\lambda\,\rho_{\rm pair}(\lambda)\,S_{\lambda}(\omega).$ (S124)
Absorbing the factor $2D_{\lambda}$ into the definition of $\rho_{\rm pair}(\lambda)$ gives
$\Pi(\omega)=\int_{0}^{\Lambda}d\lambda\,\rho_{\rm pair}(\lambda)\,\frac{1}{\lambda^{2}+\omega^{2}}.$ (S125)
This is the Cooper-channel polarization kernel quoted in the main text.
(VI) Infrared scaling of the polarization kernel: For a generic infrared TDOS
$\rho_{\rm pair}(\lambda)\sim\lambda^{\eta},\qquad\lambda\to 0,$ (S126)
Eq. (S125) gives
$\Pi(\omega)\sim\int_{0}^{\Lambda}d\lambda\,\frac{\lambda^{\eta}}{\lambda^{2}+\omega^{2}}.$ (S127)
Setting $\lambda=|\omega|x$,
$\Pi(\omega)$ $\sim|\omega|^{\eta-1}\int_{0}^{\Lambda/|\omega|}dx\,\frac{x^{\eta}}{1+x^{2}}.$ (S128)
Hence the universal infrared scaling is
$\Pi(\omega)\sim|\omega|^{\eta-1}.$ (S129)
In particular, for a flat TDOS,
$\rho_{\rm pair}(\lambda\to 0)=\rho_{0},$ (S130)
@@ Appendix D Appendix D: Derivation of the Cooper-channel polarization kernel from a relaxational mode continuum
$\Pi(\omega)$ $\simeq\rho_{0}\int_{0}^{\Lambda}\frac{d\lambda}{\lambda^{2}+\omega^{2}}$
$=\frac{\rho_{0}}{|\omega|}\tan^{-1}\!\left(\frac{\Lambda}{|\omega|}\right)$
$\xrightarrow[|\omega|\ll\Lambda]{}\frac{\pi\rho_{0}}{2}\frac{1}{|\omega|}.$ (S131)
Equation (S131) establishes the memory-dominated Cooper enhancement. A flat relaxation-rate spectrum therefore converts the marginal logarithmic pairing growth of conventional Ohmic dynamics into an algebraically enhanced infrared instability of the Cooper channel.
@@ Appendix F Appendix F: Possible connections between flat TDOS and scale–free dynamical phenomena
In this Appendix we briefly discuss a broader dynamical perspective suggested by the relaxation–spectrum framework developed in the main text. The central object of the theory is the density of relaxation rates associated with the Liouvillian generator of the coarse–grained dynamics.
Near dynamical critical regimes the spectrum of relaxation rates naturally develops a dense continuum extending toward vanishing decay rates. In particular an important limiting case is a finite infrared density of slow modes,
@@
This condition corresponds to a situation in which relaxation processes exist across arbitrarily long time scales. The system therefore cannot be characterized by a single relaxation time, and long–time memory effects emerge naturally.
(I) Long–time memory kernel: The dynamical memory kernel associated with the relaxation spectrum can be written as
@@
If the TDOS approaches a constant in the infrared, $\rho(\lambda)\approx\rho_{0}$ for $\lambda\ll\Lambda$, one obtains
@@
For long times $t\gg\Lambda^{-1}$ this yields the algebraic decay
@@
(II) Relation to scale–free relaxation statistics: Distributions of relaxation times with power–law statistics are widely encountered in complex dynamical systems, including models of self–organized criticality (SOC). In many SOC models the lifetime distribution of events takes the approximate form
@@
The relation between lifetime $\tau$ and relaxation rate $\lambda$ is
$\tau=\lambda^{-1}.$ (S153)
Since the number of modes must be invariant under a change of variables, the densities satisfy
$\rho(\lambda)\,d\lambda=P(\tau)\,d\tau.$ (S154)
Using
$\lambda=\tau^{-1},\qquad\left|\frac{d\lambda}{d\tau}\right|=\tau^{-2},$ (S155)
one obtains
$P(\tau)=\rho(\tau^{-1})\,\tau^{-2}.$ (S156)
Thus a flat infrared TDOS, $\rho(\lambda\rightarrow 0)=\rho_{0}$, implies in Eq. (S152). A scale–free lifetime distribution therefore corresponds to a relaxation spectrum that is approximately constant near $\lambda=0$.
(III) Implications for low–frequency noise: A broad distribution of relaxation times is also known to generate 1/f1/f–type noise spectra. Within the present framework the power spectrum associated with a superposition of relaxational modes takes the form
$S(\omega)\sim\int_{0}^{\Lambda}d\lambda\,\frac{\rho(\lambda)}{\lambda^{2}+\omega^{2}}.$ (S157)
Changing variables to the lifetime $\tau=1/\lambda$ and using Eq. (S156) gives
$S(\omega)\sim\int_{0}^{\infty}d\tau\,P(\tau)\,\frac{\tau^{2}}{1+(\omega\tau)^{2}}.$ (S158)
For a scale–free distribution $P(\tau)\sim\tau^{-2}$ one finds
$S(\omega)\sim\int d\tau\frac{1}{1+(\omega\tau)^{2}}.$ (S159)
Introducing the dimensionless variable $x=\omega\tau$ yields
$S(\omega)\sim\frac{1}{\omega}\int dx\,\frac{1}{1+x^{2}}.$ (S160)
Since the remaining integral is finite, the low–frequency scaling is
@@
Thus a flat TDOS naturally produces a 1/f1/f–type noise spectrum.
@@
The corresponding spectral function,
$A(\omega,q)=\frac{2\Gamma}{\left[\Delta_{q}+\rho_{0}\ln(\Lambda/|\omega|)\right]^{2}+\Gamma^{2}},$ (S163)
describes a broad continuum rather than a sharp quasiparticle peak.
@@
For $\rho(\lambda\rightarrow 0)=\rho_{0}$ one obtains
@@
which leads to an algebraically enhanced transition scale
@@
Summary. These observations suggest that the infrared organization of the relaxation spectrum may provide a common dynamical language for several scale–free phenomena observed in correlated systems. Within this perspective the central object is not a single quasiparticle mode but the entire spectrum of collective relaxation processes. Flat infrared TDOS simultaneously produces long–time memory, scale–free lifetime statistics, $1/f$ noise, and algebraic enhancement of superconducting pairing.